4. Using the above information. Fill in the remaining two columns.

Disposable Income

APC

APS

MPC

MPS

$1000

1.0

0.0

-------------

-------------

1500

0.93

.07

2000

0.90

0.10

2500

0.88

0.12

5. Based off your findings. What is the Multiplier at a disposable income of $2000?

6. If the initial spending is $10,000; what would be the total change in GDP be?
Answer #1

4. APC = Total Consumption / Disposable Income

Total Consumption = APC x Disposable Income

APS = Total Saving / Disposable Income

Total Saving = APS x Disposable Income

MPC = Change in Consumption / Change in Disposable Income

MPS = 1 - MPC

MPS = Change in Savings / Change in Disposable Income

DI 1000 1500 0.93 2000 0.9 25000.88 APC Consumption APS Savings MPC MPS 1 1000 1395 1800 2200 0 0.07 0.1 0.12 0 105 200 300 0.79 0.81 0.8 0.21 0.19 0.2

5. Multiplier = 1/MPS or 1/(1 - MPC) = 1/0.19 = 5.26

6. Change in GDP = 10,000/(1 - MPC) = 10,000/0.19 = 52631.58